# Beyond Memorization: A Rigorous Evaluation Framework for Medical Knowledge Editing

### v2

@@ A.1 Medical QA datasets
is drawn from postgraduate-level Indian medical entrance exams (AIIMS and NEET PG), spanning 2,400 healthcare topics across 21 specialties. Each question offers four answer options. We construct a subset for MedMCQAedit by selecting up to 300 QA pairs from each specialty. After filtering out entries with missing explanations, multi-correct choices, or unknown subject labels, we obtain a high-quality set used to construct MedMCQAedit.
@@
We exclude the Speech Language Pathology subset on MedExQA due to frequent inconsistencies between questions and their provided explanations. For example, one multiple-choice question lists B as the correct answer, while its accompanying explanation supports option D, rendering this subset unreliable for editing evaluations. After removing Speech Language Pathology, the remaining four specialties comprise 773 QA pairs used to construct MedExQAedit.
@@ A.2 Data Construction
We begin with two public medical QA datasets, MedExQA and MedMCQA, each of which pairs a question $q$ with a ground-truth answer $k$ and one or two human-written explanations $\mathrm{exp}$. Due to known inconsistencies (e.g., mismatches between $k$ and $\mathrm{exp}$), we filter out any QA pair whose explanation fails to support the correct answer. Concretely, we prompt an LLM (e.g., LLaMA-8B) with the $(\mathrm{exp},q)$ under in-context learning (more like in an open-book setting). If the model’s prediction does not match $k$, we discard that sample. The remaining high-quality pairs form our Verified Explanation Set $\mathcal{D}_{\mathrm{verified}}$.
@@
Next, we assess this LLM’s zero-shot performance on each question in $\mathcal{D}_{\mathrm{verified}}$ by feeding only $q$ (without $\mathrm{exp}$) in a context-free condition. Samples for which the model answers incorrectly indicate out-of-date or missing internal knowledge; these become our Original Set $\mathcal{Q}_{\mathrm{ori}}$.
@@
To generate novel datasets for rigorous Generalization and Retention assessment, we use a more powerful agent (we use DeepSeek-V3 in this study; others like GPT-4 and Gemini could also serve this role) to construct new medical questions. For each verified pair in $\mathcal{D}_{\mathrm{verified}}$, we treat its human-curated explanation that cites authoritative medical textbooks, as the medical fact. Then, the agent uses the medical fact to craft clinical-scenario QA pairs through prompting (see Figure 7), yielding candidates for $\mathcal{Q}_{\mathrm{gen}}$ and $\mathcal{Q}_{\mathrm{ret}}$.
@@
We then subject each generated variant to the same zero-shot test as in Step 2. Concretely, for generalization question candidates, we prompt the LLM without any explanation and record its predicted answer $\hat{k}$. If $\hat{k}\neq k^{\prime}$, where $k^{\prime}$ is the intended (ground-truth) answer, we include it in the generalization set $\mathcal{Q}_{\mathrm{gen}}$; As for retention question candidates, if $\hat{k}=k^{\prime}$, we add this pair to the retention set $\mathcal{Q}_{\mathrm{ret}}$. This ensures that $\mathcal{Q}_{\mathrm{gen}}$ contains only those scenarios that the LLM cannot solve without editing, while $\mathcal{Q}_{\mathrm{ret}}$ captures instances where its original knowledge remains intact.
These processes yield three sets—$\mathcal{Q}_{\mathrm{ori}}$, $\mathcal{Q}_{\mathrm{gen}}$, and $\mathcal{Q}_{\mathrm{ret}}$, which together form our final editing benchmarks MedExQAedit and MedMCQAedit.
@@ A.3 Benchmark Complexity
Our medical editing benchmarks introduce substantially more complex questions than common general-domain datasets. Whereas ZsRE questions average only 11.9 tokens, our MedMCQAedit and MedExQAedit samples exhibit mean lengths of 52.7 and 47.6 tokens for original questions, 105.8 and 100.5 for generalization questions, and 92.6 and 88.6 for retention questions, respectively (see Figure 8). This increase in question length reflects the inclusion of rich clinical context and novel scenario descriptions. In particular, $\mathcal{Q}_{\mathrm{gen}}$ and $\mathcal{Q}_{\mathrm{ret}}$ include extra patient-centered details (see Figure 9) to mimic real-world clinical scenarios, further increasing the inference challenge. By expanding both the lexical and conceptual scope of each query, our benchmarks better simulate real-world medical reasoning tasks and rigorously test an LLM’s capacity for knowledge editing under realistic complexity.
Figure 9: Examples of the three QA types. Purple text denotes the question’s domain/topic; blue text highlights the injected new correct answer; red text shows the model’s original (pre-edit) answer.
@@ Appendix B Existing Editing Methods
Fine-Tuning–Based Editing.
These traditional approaches update model parameters to incorporate new knowledge: FT+L (Zhu et al., 2020) enforces norm constraints in gradient updates to minimize interference on the unmodified facts. FT-M (Zhang et al., 2024) applies a masking strategy during training to focus updates on relevant target content. LoRA (Hu et al., 2022) introduces trainable low-rank decomposition matrices to achieve efficient adaptation with minimal additional parameters.
Parameter-Modifying Editing.
Unlike fine-tuning, these methods seek to modify only a subset of parameters to inject new facts while preserving unrelated knowledge. Broadly, they fall into two categories: i) Meta-Learning editing typically trains a hypernetwork to estimate gradient updates for knowledge insertion (De Cao et al., 2021; Tan et al., ), e.g., MEND (Mitchell et al., ) introduces a hypernetwork to transform the gradient obtained by using a low-rank decomposition to make the parameterization tractable. ii) Locate-then-Edit methods identify and then modify the weights responsible for specific factual associations: ROME (Meng et al., 2022) localizes and modifies factual associations in a transformer layer. MEMIT (Meng et al., ) extends ROME by performing batch edits across multiple critical layers for mass knowledge updates.
Recently, sequential editing methods have been introduced to support continuous updates rather than a one-off modification. PRUNE (Ma et al., 2024a) applies condition number constraints to limit perturbation to keep the general capacity of the model. AlphaEdit (Fang et al., 2024) minimizes disruption to the preserved knowledge by projecting perturbations onto the null space of its key matrices. AnyEdit (Jiang et al., 2025) decomposes long-form knowledge into sequential chunks and edits each key token autoregressive. These three newer sequential editing methods are optimized on top of the MEMIT.
@@
These methods avoid altering the parameters of the base model. Extension-based methods augment the model with external components, leaving the base parameters unchanged: GRACE (Hartvigsen et al., 2023) writes new mappings as entries in a discrete codebook in an Adaptor. SERAC (Mitchell et al., 2022) integrates a classifier and a side model to identify and learn new knowledge. Similarly, WISE (Wang et al., 2024a) deploys side FFN layers to incorporate new knowledge dynamically.
Additionally, retrieval-based methods (Song et al., 2024; Shi et al., 2024; Chen et al., 2024b) have been proposed to retrieve relevant factual information at inference time, effectively “editing” via context rather than weight changes. For example, IKE (Zheng et al., 2023) retrieves relevant factual edits and uses them to build the prompt context as input, and then prompts the model to generate updated outputs.
@@ C.1 Base Models
In this study, we evaluate editing methods on two instruction-tuned LLaMA variants: Llama-3.1-8B-Instruct and Llama-3.2-3B-Instruct  [^3] https://llama.meta.com/lama3/. We utilize layers [4, 5, 6, 7, 8] of the two LLMs for editing, except for experiments 5.3, where we investigate which layers used to edit deliver the best performance.
@@ C.2 Implementation
We implement and evaluate six representative knowledge editing methods—LoRA, ROME, MEMIT, GRACE, AnyEdit, and AlphaEdit for fair comparison. All editing workflows are built on the EasyEdit framework [^4] https://github.com/zjunlp/EasyEdit. For AnyEdit, which is not yet supported by EasyEdit, we integrate the original codebase [^5] https://github.com/jianghoucheng/AnyEdit and adopt the original hyperparameters. In terms of evaluation, we implement independent pipelines and consistently compute metrics tailored to our medical knowledge editing. During the post-edit inference phase, we employ greedy decoding to ensure deterministic outputs. Specifically, we set do_sample=False and temperature=0.0 for all evaluations, so that the edited model’s predictions reflect its learned knowledge without sampling variability. For more specific experimental settings, please see our open-source codes.
@@ C.3 Editing Paradigm
  • GTA-Edit: the ground-truth answer, with option letters stripped (e.g., remove “D:” in “D: Masson trichrome”).
  • RE-Edit: the human-written explanation excerpted from textbooks.
  • SGR-Edit: the proposed paradigm where the LLM is first prompted to generate its own chain of thought over the RE, then uses that self-generated rationale as $k^{\prime}$ for editing.
@@ C.4 Evaluation Protocol
During post-edit evaluation, we randomly permute the answer choices in each multiple-choice $\mathcal{Q}_{\mathrm{ori}}$ to prevent label memorization. All outputs follow a two-step format: rationale generation followed by the final answer. We report:
@@
Post-edit Accuracy Metrics
Efficacy $=\frac{1}{\lvert\mathcal{Q}_{\mathrm{ori}}\rvert}\sum_{i\in\mathcal{Q}_{\mathrm{ori}}}\mathbf{1}(\hat{a}_{i}=a_{i}),$ (2)
Generalization $=\frac{1}{\lvert\mathcal{Q}_{\mathrm{gen}}\rvert}\sum_{i\in\mathcal{Q}_{\mathrm{gen}}}\mathbf{1}(\hat{a}_{i}=a_{i}),$ (3)
Retention $=\frac{1}{\lvert\mathcal{Q}_{\mathrm{ret}}\rvert}\sum_{i\in\mathcal{Q}_{\mathrm{ret}}}\mathbf{1}(\hat{a}_{i}=a_{i}).$ (4)
Interpretability: ROUGE-L and BLEU scores between the injected knowledge (reference explanation or self-generated rationale) and the model’s post-edit rationale output, computed on a human-validated subset to quantify how closely the model reproduces the intended content. Detailed prompt templates, and additional examples are provided in Appendix E.
@@ Appendix D Supplementary Results and Analyses
Method Metric Pre-Edit LLaMA-8B LLaMA-3B
GTA-Edit RE-Edit SGR-Edit GTA-Edit RE-Edit SGR-Edit
LoRA Eff. 0 0.466 0.547 0.665 0.144 0.412 0.615
Gen. 0 0.416 0.460 0.503 0.337 0.439 0.492
Ret. 1 0.708 0.727 0.708 0.529 0.658 0.636
avg. 0.530 0.578 0.625 0.337 0.503 0.581
ROME Eff. 0 0.327 0.453 0.528 0.257 0.299 0.439
Gen. 0 0.296 0.403 0.447 0.251 0.406 0.385
Ret. 1 0.616 0.635 0.667 0.561 0.647 0.610
avg. 0.413 0.497 0.547 0.357 0.451 0.478
MEMIT Eff. 0 0.283 0.384 0.520 0.160 0.310 0.455
Gen. 0 0.252 0.352 0.364 0.251 0.364 0.401
Ret. 1 0.648 0.604 0.671 0.545 0.679 0.663
avg. 0.394 0.447 0.518 0.319 0.451 0.506
AnyEdit Eff. 0 0.366 0.410 0.435 0.230 0.246 0.316
Gen. 0 0.255 0.248 0.317 0.230 0.267 0.316
Ret. 1 0.789 0.758 0.764 0.781 0.747 0.754
avg. 0.470 0.472 0.505 0.414 0.420 0.462
AlphaEdit Eff. 0 0.439 0.547 0.584 0.326 0.348 0.374
Gen. 0 0.312 0.335 0.366 0.278 0.332 0.348
Ret. 1 0.867 0.795 0.789 0.775 0.743 0.754
avg. 0.539 0.559 0.580 0.460 0.474 0.492
Table 4: Complete results of GTA-Edit, RE-Edit, and SGR-Edit. Supported for RQ2-F1
D.1 Performance Comparison of Various Editing Paradigms
As shown in Table 4, relative to GTA-Edit, RE-Edit consistently raises average editing scores on LLaMA-8B by 0.2–8.4 percentage points across methods: ROME sees the largest gain (+8.4 pp), followed by MEMIT (+5.3 pp), LoRA (+4.8 pp), AlphaEdit (+2.0 pp), and AnyEdit (+0.2 pp). These improvements are driven primarily by jumps in Efficacy (up to +12.6 pp for ROME) and Generalization (up to +10.7 pp for ROME), while Retention remains stable above 61% for all methods.
Incorporating SGR-Edit yields further average gains of 2.1–7.1 pp over RE-Edit, with MEMIT (+7.1 pp) and LoRA (+4.7 pp) benefitting most. Specifically, LoRA’s combined score climbs from 57.8% to 62.5%, and MEMIT from 44.7% to 51.8%. Even AlphaEdit, which already excels under RE-Edit, improves from 55.9% to 58.0% (+2.1 pp).
Critically, these gains in Efficacy and Generalization come with only minor retention trade-offs (e.g., AnyEdit drops from 78.9% to 76.4%), confirming that richer and context-driven rationales enable deeper medical knowledge integration without undue forgetting.
Event-Driven Rationale Generation for Practical SGR-Edit

### v3

@@ A.1 Medical QA datasets
is drawn from postgraduate‐level Indian medical entrance exams (AIIMS and NEET PG), spanning 2,400 healthcare topics across 21 specialties. Each question offers four answer options (Pal et al., 2022). We construct a subset for MedMCQAedit by selecting up to 300 QA pairs from each specialty. After filtering out entries with missing explanations, multi-correct choices, or unknown subject labels, we obtain a high-quality set used to construct MedMCQAedit.
@@
We exclude the Speech Language Pathology subset on MedExQA due to frequent inconsistencies between questions and their provided explanations. For example, one multiple‐choice question lists B as the correct answer, while its accompanying explanation supports option D, rendering this subset unreliable for editing evaluations. After removing Speech Language Pathology, the remaining four specialties comprise 773 QA pairs used to construct MedExQAedit.
@@ A.2 Data Construction
We begin with two public medical QA datasets, MedExQA and MedMCQA, each of which pairs a question $q$ with a ground‐truth answer $k$ and one or two human‐written explanations $\mathrm{exp}$. Due to known inconsistencies (e.g., mismatches between $k$ and $\mathrm{exp}$), we filter out any QA pair whose explanation fails to support the correct answer. Concretely, we prompt an LLM (e.g., LLaMA-8B) with the $(\mathrm{exp},q)$ under in-context learning (more like in an open‐book setting). If the model’s prediction does not match $k$, we discard that sample. The remaining high‐quality pairs form our Verified Explanation Set $\mathcal{D}_{\mathrm{verified}}$.
@@
Next, we assess this LLM’s zero‐shot performance on each question in $\mathcal{D}_{\mathrm{verified}}$ by feeding only $q$ (without $\mathrm{exp}$) in a context‐free condition. Samples for which the model answers incorrectly indicate out-of-date or missing internal knowledge; these become our Original Set $\mathcal{Q}_{\mathrm{ori}}$.
@@
To generate novel datasets for rigorous Generalization and Retention assessment, we use a more powerful agent (we use DeepSeek-V3 in this study; others like GPT-4 and Gemini could also serve this role) to construct new medical questions. For each verified pair in $\mathcal{D}_{\mathrm{verified}}$, we treat its human‐curated explanation that cites authoritative medical textbooks, as the medical fact. Then, the agent uses the medical fact to craft clinical‐scenario QA pairs through prompting (see Figure 7), yielding candidates for $\mathcal{Q}_{\mathrm{gen}}$ and $\mathcal{Q}_{\mathrm{ret}}$.
@@
We then subject each generated variant to the same zero‐shot test as in Step 2. Concretely, for generalization question candidates, we prompt the LLM without any explanation and record its predicted answer $\hat{k}$. If $\hat{k}\neq k^{\prime}$, where $k^{\prime}$ is the intended (ground‐truth) answer, we include it in the generalization set $\mathcal{Q}_{\mathrm{gen}}$; As for retention question candidates, if $\hat{k}=k^{\prime}$, we add this pair to the retention set $\mathcal{Q}_{\mathrm{ret}}$. This ensures that $\mathcal{Q}_{\mathrm{gen}}$ contains only those scenarios that the LLM cannot solve without editing, while $\mathcal{Q}_{\mathrm{ret}}$ captures instances where its original knowledge remains intact.
[…]
@@ A.3 Benchmark Complexity
Our medical editing benchmarks introduce substantially more complex questions than common general‐domain datasets. Whereas ZsRE questions average only 11.9 tokens, our MedMCQAedit and MedExQAedit samples exhibit mean lengths of 52.7 and 47.6 tokens for original questions, 105.8 and 100.5 for generalization questions, and 92.6 and 88.6 for retention questions, respectively (see Figure 8). This increase in question length reflects the inclusion of rich clinical context and novel scenario descriptions. In particular, $\mathcal{Q}_{\mathrm{gen}}$ and $\mathcal{Q}_{\mathrm{ret}}$ include extra patient‐centered details (see Figure 9) to mimic real‐world clinical scenarios, further increasing the inference challenge. By expanding both the lexical and conceptual scope of each query, our benchmarks better simulate real‐world medical reasoning tasks and rigorously test an LLM’s capacity for knowledge editing under realistic complexity.
Steps 1 and 2 in Appendix A.2 are closely tied to how we interpret our evaluation metrics. If a base LLM were to answer all questions correctly or incorrectly, one of the evaluation sets would be empty, hindering a complete assessment of both efficacy and retention. In practice, when an LLM shows extreme pre-editing performance on a specific dataset or domain, we must dynamically adjust our focus:
  • If an LLM answers perfectly before editing, the evaluation would primarily focus on retention, which measures how well the original knowledge is retained after editing, while efficacy becomes less critical as the target questions are already answered correctly.
  • Conversely, for a poorly performing LLM, efficacy becomes the dominant metric, focusing on the model’s ability to correct its mistakes.
This adaptive interpretation ensures a meaningful assessment across all performance spectra.
While our primary objective is to provide an efficient and automated framework for knowledge editing, we suggest that for scenarios demanding exceptionally high precision in sensitive real-world deployments, introducing “human-in-the-loop” verification for an additional layer of consistency checking would indeed be a beneficial and logical extension.
Figure 9: Examples of the three QA types. Purple text denotes the question’s domain/topic; blue text highlights the injected new correct answer; red text shows the model’s original (pre‑edit) answer.
@@ Appendix B Existing Editing Methods
Fine‐Tuning–Based Editing.
These traditional approaches update model parameters to incorporate new knowledge: FT+L (Zhu et al., 2020) enforces norm constraints in gradient updates to minimize interference on the unmodified facts. FT‐M (Zhang et al., 2024) applies a masking strategy during training to focus updates on relevant target content. LoRA (Hu et al., 2022) introduces trainable low‐rank decomposition matrices to achieve efficient adaptation with minimal additional parameters.
Parameter‐Modifying Editing.
Unlike fine-tuning, these methods seek to modify only a subset of parameters to inject new facts while preserving unrelated knowledge. Broadly, they fall into two categories: i) Meta-Learning editing typically trains a hypernetwork to estimate gradient updates for knowledge insertion (De Cao et al., 2021; Tan et al., 2024), e.g., MEND (Mitchell et al., 2022a) introduces a hypernetwork to transform the gradient obtained by using a low-rank decomposition to make the parameterization tractable. ii) Locate‐then‐Edit methods identify and then modify the weights responsible for specific factual associations: ROME (Meng et al., 2022) localizes and modifies factual associations in a transformer layer. MEMIT (Meng et al., 2023) extends ROME by performing batch edits across multiple critical layers for mass knowledge updates.
Recently, sequential editing methods have been introduced to support continuous updates rather than a one-off modification. PRUNE (Ma et al., 2024a) applies condition number constraints to limit perturbation to keep the general capacity of the model. AlphaEdit (Fang et al., 2025) minimizes disruption to the preserved knowledge by projecting perturbations onto the null space of its key matrices. AnyEdit (Jiang et al., 2025) decomposes long‐form knowledge into sequential chunks and edits each key token autoregressive. These three newer sequential editing methods are optimized on top of the MEMIT.
@@
These methods avoid altering the parameters of the base model. Extension‐based methods augment the model with external components, leaving the base parameters unchanged: GRACE (Hartvigsen et al., 2023) writes new mappings as entries in a discrete codebook in an Adaptor. SERAC (Mitchell et al., 2022b) integrates a classifier and a side model to identify and learn new knowledge. Similarly, WISE (Wang et al., 2024) deploys side FFN layers to incorporate new knowledge dynamically.
Additionally, retrieval‐based methods (Song et al., 2024; Shi et al., 2024; Chen et al., 2024) have been proposed to retrieve relevant factual information at inference time, effectively “editing” via context rather than weight changes. For example, IKE (Zheng et al., 2023a) retrieves relevant factual edits and uses them to build the prompt context as input, and then prompts the model to generate updated outputs.
@@ C.1 Base Models
In this study, we evaluate editing methods on two instruction‐tuned LLaMA variants: Llama-3.1-8B-Instruct and Llama-3.2-3B-Instruct  [^3] https://llama.meta.com/lama3/. We utilize layers [4, 5, 6, 7, 8] for editing based on the findings from Appendix D.1.
@@ C.2 Implementation
We implement and evaluate six representative knowledge editing methods—LoRA, ROME, MEMIT, GRACE, AnyEdit, and AlphaEdit for fair comparison. All editing workflows are built on the EasyEdit framework [^4] https://github.com/zjunlp/EasyEdit. For AnyEdit, which is not yet supported by EasyEdit, we integrate the original codebase [^5] https://github.com/jianghoucheng/AnyEdit and adopt the original hyperparameters. In terms of evaluation, we implement independent pipelines and consistently compute metrics tailored to our medical knowledge editing. During the post‐edit inference phase, we employ greedy decoding to ensure deterministic outputs. Specifically, we set do_sample=False and temperature=0.0 for all evaluations, so that the edited model’s predictions reflect its learned knowledge without sampling variability. For locate-and-edit KE methods, we select layers 4-8 as the target knowledge editing layers for all subsequent experiments (see Appendix D.1 for a detailed analysis). For more specific experimental settings, please see our open-source codes.
@@ C.3 Editing Paradigm
  • GTA-Edit: model is edited through a ground‐truth answer, with option letters stripped (e.g., remove “D:” in “D: Masson trichrome”).
  • RE-Edit: editing the model with the human‐written explanation excerpted from textbooks.
  • SGR-Edit: the proposed paradigm where the LLM is first prompted to generate its own chain of thought over the RE, then uses that self‐generated rationale as $k^{\prime}$ for editing. The prompt used for rationale generation is in Figure˜10.
@@ C.4 Evaluation Protocol
During post‐edit evaluation, we randomly permute the answer choices in each multiple‐choice $\mathcal{Q}_{\mathrm{ori}}$ to prevent label memorization. All outputs follow a two‐step format: rationale generation followed by the final answer. We report:
@@
Post‐edit Accuracy Metrics
Efficacy $=\frac{1}{\lvert\mathcal{Q}_{\mathrm{ori}}\rvert}\sum_{i\in\mathcal{Q}_{\mathrm{ori}}}\textbf{1}(\hat{a}_{i}=a_{i}),$ (2)
Generalization $=\frac{1}{\lvert\mathcal{Q}_{\mathrm{gen}}\rvert}\sum_{i\in\mathcal{Q}_{\mathrm{gen}}}\textbf{1}(\hat{a}_{i}=a_{i}),$ (3)
Retention $=\frac{1}{\lvert\mathcal{Q}_{\mathrm{ret}}\rvert}\sum_{i\in\mathcal{Q}_{\mathrm{ret}}}\textbf{1}(\hat{a}_{i}=a_{i}).$ (4)
Interpretability: ROUGE-L and BLEU scores between the injected knowledge (reference explanation or self-generated rationale) and the model’s post-edit rationale output, computed on a human-validated subset to quantify how closely the model reproduces the intended content. Furthermore, we adopt Quality of Reasoning (QoR) as a normalized (0–1) score via an AI judge (DeepSeek-V3), assessing the generated rationales across five dimensions: factual accuracy, logical flow, relevance, completeness, and answer correctness (see prompt in Figure˜14) Additional examples are provided in Appendix E.
Figure 11: LLM-as-a-judge prompt for Rationale Factuality and Hallucination Risk evaluations
@@ Appendix D Supplementary Results and Analyses
D.1 Localization of Medical Knowledge in LLMs
Figure 12: Layer‐Wise editing performance
This section investigates the layer-wise storage of medical knowledge in LLMs. We edit knowledge across four disjoint layers to check the storage of certain knowledge: For LLaMA-8B, we select layers 4–8, 11–15, 18–22, and 25–29; for LLaMA-3B and Qwen-7B, we use layers 4–8, 10–14, 16–20, and 22–26. We sample 100 QA pairs from MedMCQAedit and group target knowledge by token length to assess the impact of knowledge granularity. GTA targets are the shortest answers (<10 tokens), RE explanations span 50–100 or 100–150 tokens, and SGR rationales range from 150–200 to 200–250 tokens. Based on Figure 12, we draw the following findings:
Medical knowledge is primarily stored in shallower layers.
For both LLaMA-8B and 3B, editing operations targeting shallower layers 4-8 consistently yield the highest effectiveness. When editing on LLaMA-8B, average scores in layers 4-8 range from 0.433 to 0.583, and for LLaMA-3B, they span from 0.333 to 0.533, consistently outperforming edits in deeper layers. Despite Qwen-7B also achieving strong editing effectiveness in deep layers (all above 40%), the most effective network layers for it still reside in the shallower layers.
Across all editing paradigms, the most effective edits are consistently achieved in the shallower layers.
Based on the above experiments and analysis, we adopt the shallower layers 4-8 as the knowledge editing layers for our experiments.
D.2 Factual Grounding and Rationale Verification
Subset Samples Avg. Factual Halluc. Halluc.
Score Count Rate
MedExQAedit (LLaMA-3B) 60 4.93 1 1.67%
MedExQAedit (LLaMA-8B) 46 5.00 0 0.00%
MedExQAedit (Qwen-7B) 50 4.90 2 4.00%
MedMCQAedit (LLaMA-3B) 100 4.91 3 3.00%
MedMCQAedit (LLaMA-8B) 100 5.00 0 0.00%
MedMCQAedit (Qwen-7B) 100 4.99 0 0.00%
Table 4: Factual evaluation and hallucination analysis across different subsets and models. The prompt is provided in Figure 11
To ensure the factual grounding of these rationales for both benchmark construction and the proposed editing paradigm, we performed a targeted evaluation on a subset to assess rational factuality. Our methodology involved two stages:
  1. 1.
    AI Judge Evaluation: We first utilized DeepSeek-V3 as an AI judge to evaluate the generated rationales (Zheng et al., 2023b). DeepSeek-V3 assessed the rationales across several dimensions, including Factual Alignment, Hallucination, Logical Consistency, Completeness, and Accuracy, ultimately assigning a factual score on a 1-5 scale and indicating whether hallucination was detected.
  2. 2.
    Human Verification: Recognizing that even an advanced AI judge might exhibit its own form of hallucination in knowledge-specific tasks (Szymanski et al., 2025), especially with complex medical knowledge, we further introduced three human verifiers. These human annotators meticulously reviewed a subset of the rationales to confirm their factual accuracy and presence of hallucination, validating the AI judge’s assessment.
As observed from Table 4, the average factual scores are remarkably high (ranging from 4.91 to 5.0), indicating strong factual alignment and consistency. Crucially, the hallucination rates are extremely low, with our rigorous verification process (involving both the AI judge and human annotators) confirming the high factual consistency of the generated rationales. This highlights that LLMs can self-generate high-quality rationales to update their knowledge.
D.3 Event‐Driven Rationale Generation for Practical SGR‐Edit